Management Meeting

2019-12-13 12:00
America/New York
Friday, 13 December 2019
, at 17:00 (GMT), duration : 00:00

Agenda:
  • Run report
    • Collisions now underway. Intensity (as of Wed) still a little low. 14 min fills likely operation mode.
    • TPC Lasers were fixed
    • eTOF data now in QA plots. Still some issues with DAQ/T0 being investigated. Stability of operations ongoing, progress mass
    • Other detectors coming up well
    • Weekly QA meeting started. New tool almost ready to automate the QA plots for the PWG - a summer project of Ben and an summer high school student at BNL. HLT and Fast Offline ongoing.
    • QA fast of Shift need to start a week earlier than planned. This is happening and experienced folks identified to take on the task as it ramps up.
  •  Production
    • Run 18: FxT. ToF calibration ongoing.
    • Run 19: TPC timing inner vs outer
    • Run 17: Production calibration samples. Look better but studies still ongoing, not yet perfect. Need to understand how good it needs to be for analyses planned.
  • Embedding
    • Run 16-18 TPC simulation of clusters not accurate enough due to ADC threshold change. Work ongoing.
    • Run 15 pp200. Data and simulation don't match. Material from HFT seems to be correct. Rather maybe GEANT run incorrectly. Being investigated.
    • Large net-proton embedding request submitted for 27 GeV. This is the PAC request to check non-binomial correction is correct